# Why Reinforcement Fine-Tuning Enables MLLMs Preserve Prior Knowledge Better: A Data Perspective

### v3

@@ 5.4 Learning Dynamics-based analysis of Data Distribution
Following ren2024learning, we employ learning dynamics to describe “how the change in parameter induced by a step of gradient descent on single training example impacts the probability of another example ”. Here, we treat as a GRPO-rollout sample or man-made SFT sample, and as a sample of prior knowledge. We have
@@
Proof of the theorem and more discussion can be found in Appendix D. The theorem shows that the effect of on is mainly determined by three factors: (1) the model’s sensitivity to the old and new knowledge ( and ), and (2) the level of interference between them, captured by . Since the gradients with respect to the logits (i.e., and ) are typically bounded, this implies that the relative interference is the dominant factor driving forgetting. A larger means more interference between and . Besides, our analysis in this section also depends on the assumption of “the eNTK matrix remains roughly stable over training”, which is well-validated in ren2024learning and our following experiments.
@@ 5.5 What makes the model-generated Reasoning Data different?
According to Theorem 5.2 (Proof in Appendix D), we find that the influence of learning on is nearly the same as the influence of learning on . Additionally, since the eNTK matrix stabilizes (ren2024learning) in the later stages of pretraining, the interactions between and remain consistent across the training step , also observed in Fig. 2. Therefore, models pretrained with prior knowledge show lower perplexity for Rea-GRPO-Rollout, indicating that training with prior knowledge enhances these samples. During post-training, further training on the Rea-GRPO-Rollout samples results in less interference with prior knowledge compared to other higher perplexity samples like Rea-4o-Rollout. As shown in Appendix Fig. 8, when training with Rea-GRPO-Rollout, perplexity of sentences representing prior knowledge continues to decrease on 3B model and remains low on 7B model. In contrast, under Rea-4o-Rollout, perplexity steadily increases. As a result, forgetting effect for Rea-GRPO-Rollout is less pronounced than for Rea-4o-Rollout. Moreover, reinforcement learning algorithms like GRPO, which naturally generate training samples through model rollouts, tend to produce samples with lower perplexity under the base model. This explains why reinforcement learning methods are less prone to catastrophic forgetting.
@@ 5.7 LLM Experiments on Math Reasoning and ScienceQA
We additionally provide experiments on the LLM Qwen2.5-Instruct (qwen2.5) here, showing its forgetting behavior during post-training on math reasoning, along with the corresponding results. We hope these extra experiments can further strengthen the generality and credibility of our theoretical analysis and conclusions. More detailed experiments setup can be found in Appendix F. As summarized in Tab. 2, the math reasoning experiments exhibit a forgetting pattern highly consistent with our multimodal jigsaw setting: on both 3B and 7B scales, SFT-Non-Rea achieves the largest performance drop on the old math (GSM8K, MATH-500) and instruction-following (IFEval) benchmarks, while reasoning-augmented SFT with external CoT (SFT-Rea-4o-Rollout) forgets less but still substantially more than SFT-Rea-GRPO-Rollout. The latter attains strong gains on the new ORZ task while keeping the performance on old tasks close to the base models, indicating the same hierarchy of forgetting severity, i.e., Non-Rea Rea-4o Rea-GRPO.
@@
Beyond math reasoning, we also conduct experiments on a scientific multiple-choice QA benchmark to test the robustness of our conclusions. Concretely, we use the Sci-MCQ4 subset from SciKnowEval (sciknoweval); detailed settings and full results are provided in Appendix H. As shown in Tab. 9 and Fig. 16, 17, we again observe the same hierarchy of forgetting severity on prior benchmarks, while SFT-Rea-GRPO-Rollout achieves the best trade-off between performance gains on the new Sci-MCQ4 task and retention on old tasks, further supporting the generality of our analysis.
| Qwen2.5-3B-Instruct | Qwen2.5-7B-Instruct | |||||||||||||||||
| Base | RFT | SFT-Non-Rea |
Base | RFT | SFT-Non-Rea |
| Open-Reasoner-Zero (test) (New Task) | ||||||||||||||||||
| Training steps | – | 2,650 | 1,600 | 2,140 | 2,140 | – | 2,650 | 1,600 | 2,140 | 2,140 | ||||||||
| ORZ Test | 21.3 | 35.0 (13.7) | 23.4 (2.1) | 35.4 (14.0) | 37.8 (16.4) | 32.1 | 49.3 (17.2) | 30.3 (1.8) | 45.0 (12.9) | 53.4 (21.3) | ||||||||
| Math Reasoning (Old Tasks) | ||||||||||||||||||
| GSM8k | 84.1 | 83.4 (0.7) | 15.1 (69.0) | 79.9 (9.2) | 83.0 (1.1) | 90.1 | 90.2 (0.1) | 21.8 (68.4) | 85.8 (4.3) | 90.3 (0.2) | ||||||||
| Math-500 | 42.4 | 55.2 (12.8) | 19.4 (23) | 50.8 (8.4) | 54.4 (12.0) | 66.6 | 64.8 (1.8) | 26.4 (40.2) | 57.2 (9.4) | 66.4 (0.2) | ||||||||
| Instruction Following (Old Task) | ||||||||||||||||||
| IFEval | 71.6 | 73.4 (1.8) | 64.0 (7.6) | 68.0 (3.6) | 72.7 (1.1) | 80.6 | 80.5 (0.1) | 57.2 (23.4) | 64.4 (16.2) | 80.0 (0.6) | ||||||||
@@ 6 Discussion and Conclusion
SFT is a widely used post-training method and is often employed as a cold-start phase for RFT (DeepSeek-R1), helping the model acquire basic skills that support subsequent exploration. Besides, SFT also enables the base model to master novel tasks quickly. However, manually curated SFT corpora can lead to the forgetting of prior knowledge. In this work, we show that one can instead construct more stable SFT training data from the model’s own reasoning trajectories produced by RFT. Even a short RFT phase is sufficient to generate such self-consistent data (Sec. 5.8), and a subsequent SFT update on this corpus attains new-task performance comparable to RFT while preserving prior knowledge better than SFT-Rea-4o-Rollout. Therefore, developing an efficient and reliable interplay between SFT and RFT that combines their respective advantages remains a promising problem.
@@ Appendix C Connection between GRPO and SFT
This section follows the discussion of DeepSeekMath (shao2024deepseekmath) on the unified paradigm of GRPO and SFT closely. And we include the derivation here for the completeness of the paper. We will first derive the gradient of GRPO loss. Specifically, we use the following unbiased estimator as our KL divergence loss:
@@ Appendix D Proof of Learning Dynamics Related Theorem
The second term in this decomposition, , is called the empirical neural tangent kernel and can evolve during training as the network adapts. For sufficiently wide networks initialized properly and trained with small learning rates, stays nearly fixed throughout training—the limiting kernel in this case is referred to as the neural tangent kernel (arora2019exact; jacot2018neural; ren2024learning). Additionally, ren2024learning also validated a relaxed assumption for LLM fine-tuning: the relative influence of learning on other inputs remains roughly stable over training. Besides, the optimization steps during post-training of MLLMs in our paper are very less compared to the steps used in pre-training. So, the relative influence between and during the post-training remains similar to the influence during pre-training is a reasonable hypothesis.
@@ Appendix E More Results of Jigsaw Puzzles
GRPO with Periodic Restart. We observe that training with GRPO for too many epochs on a novel task can lead to instability. Periodic restarts are essential for maintaining stable training. As shown in Fig. 9, prolonged training eventually causes the KL divergence loss to explode, consistent with findings from ProRL. To address this, we reset both the optimizer state and the reference policy every 2 epochs. Accordingly, when rule-based constraints or well-shaped reward signals are available, a prolonged exploration stage helps the model generate higher-quality trajectories. This allows it to acquire new skills that are not present in the base model, leading to measurable improvements on the target tasks.
@@ Appendix F More Details of LLM on Math Reasoning
Math Reasoning Dataset. We use the curated math corpus released by Open-Reasoner-Zero (ORZ) as our large-scale reasoning-oriented training data, and randomly split it into training and held-out test data. We refer to this held-out split as ORZ Test, which serves as a new target task for post-training. Each example is a competition-style math problem paired with a verifiable final answer, without any visual input.
LLMs and Evaluation. For math experiments, we use Qwen2.5-3B-Instruct (qwen2.5) and Qwen2.5-7B-Instruct as base LLMs. We treat ORZ Test as a new target task and report answer accuracy on it, and use GSM8K (gsm8k), MATH-500 (MATH-500), and IFEval (IFEval) as prior knowledge to monitor retention of prior math and instruction-following abilities, with their standard accuracy metrics.
@@ Appendix H LLM Experiments on Scientific Multiple-Choice QA
Scientific MCQ Dataset. We further investigate catastrophic forgetting on scientific multiple-choice questions using the Sci-MCQ4 subset from SciKnowEval (sciknoweval). From the original corpus, we randomly sample 8,500 examples and split them into training and held-out test data. Each instance is a four-choice science question together with its correct option, covering multi-level scientific knowledge such as physics, chemistry, and biology. We refer to this held-out split as Sci-MCQ4 Test, which serves as a target task for post-training.

### v4

@@ 5.4 Learning Dynamics-based analysis of Data Distribution
Following Ren and Sutherland (2024), we employ learning dynamics to describe “how the change in parameter induced by a step of gradient descent on single training example impacts the probability of another example ”. Here, we treat as a GRPO-rollout sample or man-made SFT sample, and as a sample of prior knowledge. We have
@@
Proof of the theorem and more discussion can be found in Appendix D. The theorem shows that the effect of on is mainly determined by three factors: (1) the model’s sensitivity to the old and new knowledge ( and ), and (2) the level of interference between them, captured by . Since the gradients with respect to the logits (i.e., and ) are typically bounded, this implies that the relative interference is the dominant factor driving forgetting. A larger means more interference between and . Besides, our analysis in this section also depends on the assumption of “the eNTK matrix remains roughly stable over training”, which is well-validated in Ren and Sutherland (2024) and our following experiments.
@@ 5.5 What makes the model-generated Reasoning Data different?
According to Theorem 5.2 (Proof in Appendix D), we find that the influence of learning on is nearly the same as the influence of learning on . Additionally, since the eNTK matrix stabilizes (Ren and Sutherland, 2024) in the later stages of pretraining, the interactions between and remain consistent across the training step , also observed in Fig. 2. Therefore, models pretrained with prior knowledge show lower perplexity for Rea-GRPO-Rollout, indicating that training with prior knowledge enhances these samples. During post-training, further training on the Rea-GRPO-Rollout samples results in less interference with prior knowledge compared to other higher perplexity samples like Rea-4o-Rollout. As shown in Appendix Fig. 8, when training with Rea-GRPO-Rollout, perplexity of sentences representing prior knowledge continues to decrease on 3B model and remains low on 7B model. In contrast, under Rea-4o-Rollout, perplexity steadily increases. As a result, forgetting effect for Rea-GRPO-Rollout is less pronounced than for Rea-4o-Rollout. Moreover, reinforcement learning algorithms like GRPO, which naturally generate training samples through model rollouts, tend to produce samples with lower perplexity under the base model. This explains why reinforcement learning methods are less prone to catastrophic forgetting.
@@ 5.7 LLM Experiments on Math Reasoning and ScienceQA
We additionally provide experiments on the LLM Qwen2.5-Instruct (Yang et al., 2024) here, showing its forgetting behavior during post-training on math reasoning, along with the corresponding results. We hope these extra experiments can further strengthen the generality and credibility of our theoretical analysis and conclusions. More detailed experiments setup can be found in Appendix F. As summarized in Tab. 2, the math reasoning experiments exhibit a forgetting pattern highly consistent with our multimodal jigsaw setting: on both 3B and 7B scales, SFT-Non-Rea achieves the largest performance drop on the old math (GSM8K, MATH-500) and instruction-following (IFEval) benchmarks, while reasoning-augmented SFT with external CoT (SFT-Rea-4o-Rollout) forgets less but still substantially more than SFT-Rea-GRPO-Rollout. The latter attains strong gains on the new ORZ task while keeping the performance on old tasks close to the base models, indicating the same hierarchy of forgetting severity, i.e., Non-Rea Rea-4o Rea-GRPO.
@@
Beyond math reasoning, we also conduct experiments on a scientific multiple-choice QA benchmark to test the robustness of our conclusions. Concretely, we use the Sci-MCQ4 subset from SciKnowEval (Feng et al., 2024); detailed settings and full results are provided in Appendix H. As shown in Tab. 9 and Fig. 16, 17, we again observe the same hierarchy of forgetting severity on prior benchmarks, while SFT-Rea-GRPO-Rollout achieves the best trade-off between performance gains on the new Sci-MCQ4 task and retention on old tasks, further supporting the generality of our analysis.
| Qwen2.5-3B-Instruct | Qwen2.5-7B-Instruct | |||||||||
| Base | RFT | SFT-Non-Rea | SFT-Rea- 4o-Rollout | SFT-Rea- GRPO-Rollout | Base | RFT | SFT-Non-Rea | SFT-Rea- 4o-Rollout | SFT-Rea- GRPO-Rollout | |
| Open-Reasoner-Zero (test) (New Task) | ||||||||||
| Training steps | – | 2,650 | 1,600 | 2,140 | 2,140 | – | 2,650 | 1,600 | 2,140 | 2,140 |
| ORZ Test | 21.3 | 35.0 (13.7) | 23.4 (2.1) | 35.4 (14.0) | 37.8 (16.4) | 32.1 | 49.3 (17.2) | 30.3 (1.8) | 45.0 (12.9) | 53.4 (21.3) |
| Math Reasoning (Old Tasks) | ||||||||||
| GSM8k | 84.1 | 83.4 (0.7) | 15.1 (69.0) | 79.9 (9.2) | 83.0 (1.1) | 90.1 | 90.2 (0.1) | 21.8 (68.4) | 85.8 (4.3) | 90.3 (0.2) |
| Math-500 | 42.4 | 55.2 (12.8) | 19.4 (23) | 50.8 (8.4) | 54.4 (12.0) | 66.6 | 64.8 (1.8) | 26.4 (40.2) | 57.2 (9.4) | 66.4 (0.2) |
| Instruction Following (Old Task) | ||||||||||
| IFEval | 71.6 | 73.4 (1.8) | 64.0 (7.6) | 68.0 (3.6) | 72.7 (1.1) | 80.6 | 80.5 (0.1) | 57.2 (23.4) | 64.4 (16.2) | 80.0 (0.6) |
@@ 6 Discussion and Conclusion
SFT is a widely used post-training method and is often employed as a cold-start phase for RFT (DeepSeek-AI et al., 2025), helping the model acquire basic skills that support subsequent exploration. Besides, SFT also enables the base model to master novel tasks quickly. However, manually curated SFT corpora can lead to the forgetting of prior knowledge. In this work, we show that one can instead construct more stable SFT training data from the model’s own reasoning trajectories produced by RFT. Even a short RFT phase is sufficient to generate such self-consistent data (Sec. 5.8), and a subsequent SFT update on this corpus attains new-task performance comparable to RFT while preserving prior knowledge better than SFT-Rea-4o-Rollout. Therefore, developing an efficient and reliable interplay between SFT and RFT that combines their respective advantages remains a promising problem.
@@
Acknowledgments
The authors wish to thank the anonymous reviewers for their helpful comments. This work was partially funded by Henan Province Major Industrial “Challenge-Based Innovation” (No. 251000210300), National Natural Science Foundation of China (No.62476061, 62376061, 62576106, 62521004), Shanghai Rising-Star Program (23QA1400200), and Natural Science Foundation of Shanghai (23ZR1403500). Supported by Shanghai Artificial Intelligence Laboratory.
@@ Appendix C Connection between GRPO and SFT
This section follows the discussion of DeepSeekMath (Shao et al., 2024) on the unified paradigm of GRPO and SFT closely. And we include the derivation here for the completeness of the paper. We will first derive the gradient of GRPO loss. Specifically, we use the following unbiased estimator as our KL divergence loss:
@@ Appendix D Proof of Learning Dynamics Related Theorem
The second term in this decomposition, , is called the empirical neural tangent kernel and can evolve during training as the network adapts. For sufficiently wide networks initialized properly and trained with small learning rates, stays nearly fixed throughout training—the limiting kernel in this case is referred to as the neural tangent kernel (Arora et al., 2019; Jacot et al., 2018; Ren and Sutherland, 2024). Additionally, Ren and Sutherland (2024) also validated a relaxed assumption for LLM fine-tuning: the relative influence of learning on other inputs remains roughly stable over training. Besides, the optimization steps during post-training of MLLMs in our paper are very less compared to the steps used in pre-training. So, the relative influence between and during the post-training remains similar to the influence during pre-training is a reasonable hypothesis.
@@ Appendix E More Results of Jigsaw Puzzles
GRPO with Periodic Restart. We observe that training with GRPO for too many epochs on a novel task can lead to instability. Periodic restarts are essential for maintaining stable training. As shown in Fig. 9, prolonged training eventually causes the KL divergence loss to explode, consistent with findings from Liu et al. (2025a). To address this, we reset both the optimizer state and the reference policy every 2 epochs. Accordingly, when rule-based constraints or well-shaped reward signals are available, a prolonged exploration stage helps the model generate higher-quality trajectories. This allows it to acquire new skills that are not present in the base model, leading to measurable improvements on the target tasks.
@@ Appendix F More Details of LLM on Math Reasoning
Math Reasoning Dataset. We use the curated math corpus released by Open-Reasoner-Zero (Hu et al., 2025) as our large-scale reasoning-oriented training data, and randomly split it into training and held-out test data. We refer to this held-out split as ORZ Test, which serves as a new target task for post-training. Each example is a competition-style math problem paired with a verifiable final answer, without any visual input.
LLMs and Evaluation. For math experiments, we use Qwen2.5-3B-Instruct (Yang et al., 2024) and Qwen2.5-7B-Instruct as base LLMs. We treat ORZ Test as a new target task and report answer accuracy on it, and use GSM8K (Cobbe et al., 2021), MATH-500 (Hendrycks et al., 2021), and IFEval (Zhou et al., 2023) as prior knowledge to monitor retention of prior math and instruction-following abilities, with their standard accuracy metrics.
@@ Appendix H LLM Experiments on Scientific Multiple-Choice QA
Scientific MCQ Dataset. We further investigate catastrophic forgetting on scientific multiple-choice questions using the Sci-MCQ4 subset from SciKnowEval (Feng et al., 2024). From the original corpus, we randomly sample 8,500 examples and split them into training and held-out test data. Each instance is a four-choice science question together with its correct option, covering multi-level scientific knowledge such as physics, chemistry, and biology. We refer to this held-out split as Sci-MCQ4 Test, which serves as a target task for post-training.